OpenStudy (asib1214):

what volume of hydrogen gas at STP will be produced when 1000g of aluminum is added to 4.00L of 1.40mol/L Sulfuric acid? @abb0t

OpenStudy (asib1214):

@aaronq

OpenStudy (aaronq):

so you got the balanced equation?

OpenStudy (asib1214):

2Al + 6HCl ---> 2AlCl3 + 3H2

OpenStudy (aaronq):

okay now, find the limiting reagent. we divide the moles of each of the reactants by their stoichiometric coefficient and see which one has less.

OpenStudy (asib1214):

ok please wait let me do that step

OpenStudy (aaronq):

okay

OpenStudy (asib1214):

sorry it's not coming to mind :( please give me a clue :(

OpenStudy (aaronq):

okay, so for the aluminum, you wanna find the moles from the mass. \(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (asib1214):

yah but we is the mass?

OpenStudy (aaronq):

For the sulfuric acid, you need the moles from the molarity and the volume \(\sf Molarity=\dfrac{moles}{L}\)

OpenStudy (asib1214):

we don't have the mass given in this question

OpenStudy (aaronq):

1000g of aluminum

OpenStudy (asib1214):

ok thanks...

OpenStudy (asib1214):

so would it be 1000g/36.46 = 27.42mol

OpenStudy (asib1214):

of cl2

OpenStudy (aaronq):

now find the moles of sulfuric acid

OpenStudy (aaronq):

that should've been aluminium

OpenStudy (asib1214):

oh sorry...

OpenStudy (asib1214):

yah aluminum

OpenStudy (asib1214):

yah but i don't have the mount of Cl2

OpenStudy (asib1214):

in grams

OpenStudy (aaronq):

It's not even asking for \(Cl_2\)

OpenStudy (aaronq):

it's not in the question

OpenStudy (asib1214):

sorry its 5.6 mol

OpenStudy (asib1214):

sorry i mean sulfuric acid :(

OpenStudy (asib1214):

i'm having a really bad day :'(

OpenStudy (aaronq):

okay, divide each by their coefficients

OpenStudy (aaronq):

aw sorry to hear that

OpenStudy (asib1214):